If f(x) = sin¹x, show that(1-x2)f"(x) xf'(x)

1
Expert's answer
2022-05-03T03:00:42-0400
"f(x)=\sin ^{-1}x"

"f'(x)=\dfrac{1}{\sqrt{1-x^2}}"

"f''(x)=-\dfrac{1}{2(\sqrt{1-x^2})^3}(-2x)=\dfrac{x}{(1-x^2)^{3/2}}"

"(1-x^2)f''(x)-xf'(x)"

"=(1-x^2)(\dfrac{x}{(1-x^2)^{3/2}})-x(\dfrac{1}{\sqrt{1-x^2}})"

"=\dfrac{x}{\sqrt{1-x^2}}-\dfrac{x}{\sqrt{1-x^2}}=0"